"\intop" 1/x2-4x+3 dx=?


1
Expert's answer
2022-01-12T18:26:40-0500

By the sum rule,

"\int(\frac{1}{x²}-4x+3)dx="

"\int \frac{1}{x²}dx-\int4xdx+\int3dx"

Integrating each integral, we have

"=-\frac{1}{x}-2x²+3x+c"

( "c" is added since it's an indefinite integral).